OpenStudy (anonymous):

find the midpoint of the segments with the given endpoints. a(-4 -7) and c(4 4)

OpenStudy (anonymous):

\[\left( \frac{ x _{1}+x _{2} }{ 2 },\frac{ y _{1}+y _{2} }{ 2 }\right)\]

OpenStudy (cggurumanjunath):